A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC.
1
Expert's answer
2020-05-08T15:27:46-0400

"R_t=R_0(1+\alpha t)"


For carbon "\alpha =-5\cdot10^{-4}°C^{-1}" at "0\u00b0C".


If "R_0=500 \Omega" at "t=0\u00b0C".


So, we have


"R_t=R_0(1+\alpha t)=500(1-0.0005\cdot50)=487.5 \Omega"